20130217 Laserwire PLIC Detector
As a possible way of measuring background a secondary PLIC detector was made.
This was done quite quickly and there wasn't sufficient time to develop it fully. A ring made of bits of plastic pipe was made to coil the optical fibre through and around but this took too long so the rest was simply curled up and strapped to the beam pipe. It was placed between QM10AFF and QM10BFF beside the REF3 cavity. The fibre ends are put through a rubber stopper directly into the PMT and taped to prevent any light coming in.
- At 1250V, there were no observable dark current pulses on the scope.
- At 1250V, and current ATF2 optics, the signal was approximately 100ns long and between 200mV and 1.5V (almost saturation) depending on the beam condition.
Location
Arrows show location of fibre coil and PMT
Close Up
High Voltage Supply
The high voltage is supplied from the laserwire hut (red power supply in 3rd rack, at the bottom).
- Select channel 2 using up and down arrows.
- Check black and white manual dials are set to 1250V in the middle of the power supply
- Press 'L' to load the voltage
- Press 'Z' to zero the voltage on that channel
Signal
The signal is currently plugged into the scope - the scope currently has two inputs, one is a signal from the BPM at the laserwire IP, the other is this PLIC detector. It has BNC 50 ohm termination and has a green sleeve on the cable.
The laser-wire camac has two 16 channel ADCs. One is used for the laser-wire signal and other for the laser photodiode before the laserwire IP. These are put into two separate ADCs as they require different gates.
Should you want to use this detector, it will need to be attenuated (bottom left of middle rack), gated (blue gate generator middle rack, bottom, middle of nim bin) and digitised (need to sacrifice either laserwire signal or photodiode signal for this).
Signal & High Voltage Supply